Can someone help me solve this. 3/4x-2=5+2/3x
Answer by stanbon(75887)

(3/4)x-2=5+(2/3)x
Add 2 to both sides to get:
(3/4)x = 7 + (2/3)x
Multiply thru by 12 to get:
9x = 84 + 8x
x = 84
--------------
Cheers,
Stan H.